Salika travels to school by train every day. The probability that her train will be late on any day is 0.3. If salika is late on 4 consecutive days she gets a detention. what is the probability she will get a detention during a week?

we would draw a tree diagram to show this.
4 consecutive days means everyday except friday, or everyday except monday.
this would be 2(0.3x0.3x0.3x0.3x0.7)

EC

Related Maths GCSE answers

All answers ▸

Find the length of a side of the triangle (Pythagoras' Theorem) Two sides are of length 3cm and 4cm, find the length of the Hypotenuse.


(a) show that 3/10 + 2/15 = 13/30 (b) show that 2 5/8 ÷ 1 1/6 = 2 1/4


The circle c has equation x^2 + y^2 = 1. The line l has gradient 3 and intercepts the y axis at the point (0, 1). c and l intersect at two points. Find the co-ordinates of these points.


Make x the subject of the formula 3(2x – y) = ax – 4